OpenStudy (goformit100):

Hasan buys two kinds of cloth materials for school uniforms, shirt material that costs him Rs 50 per metre and trouser material that costs him Rs 90 per metre. For every 2 meters of the trouser material he buys 3 metres of the shirt material. He sells the materials at 12% and 10% profit respectively. His total sale is Rs 36660. How much trouser material did he buy?

OpenStudy (justin9529):

shirt material = Rs 50 per metre profit=12% sp =(50*12/100) +50=56 trouser material = Rs 90 per metre. profit=10% sp =(90*10/100) +90=99 For every 2 meters of the trouser material he buys 3 metres of the shirt material. = (56*3)+(99*20) =366 let the number of sets he sold be y.

OpenStudy (goformit100):

Ok, after this ?

OpenStudy (justin9529):

when you let the number of sets be y: total sell = 366*y=36600 so,y=100 every set contains 2 meters of trouser material. so Hasan sold 2*100=200 meters.

OpenStudy (goformit100):

Thankyou Sir

OpenStudy (justin9529):

Welcome:)